Please help me with this question. I really need step by step. Thanks so much.
1/6x - 1/8x = 1/72

Answer by checkley77(12569)

1/6x - 1/8x = 1/72
x/6-x/8=1/72
(8x-6x)/48=1/72
2x/48=1/72
2x*72=1*48
144x=48
x=48/144
x=1/3 ans.
Proof:
(1/3)/6-(1/3)/8=1/72
1.3*1/6-1/3*1/8=1/72
1/18-1/24=1/72
(24-18)/(18*24)=1/72
6/432=1/72
1/72=1/72